# Effects of Mosaic Crystal Instrument Functions on X-ray Thomson Scattering Diagnostics

### v1

@@ III.3 Source-and-Instrument Function as a Convolution
In Fig. 10, we compare calculations of the XRTS spectrum using the convolution and for a full kernel calculation. To estimate the energy dependence of the mosaicity and IRC, we use linear fits to the data reported Ref. Gerlach et al. (2015), maintaining the gradients but shifting the values to match the mosaic spread and IRC widths we report in Table 1 (b).
@@
In backwards scattering, the differences between the full kernel calculation and convolution become more apparent, with the relative difference increasing further from the quasi-elastic feature. This is because, compared to forward scattering, there is substantial signal at energy losses far from the quasi-elastic feature in backwards scattering which will contribute to the differences in shape across the spectrum. Additionally, there is a notable notable drop in intensity above 8.6 keV – this is because the crystal position was set only to measure up to $\sim 100$ eV above the elastic (cf. Fig. 7 (b)), so it is perhaps not surprising to see substantial differences between the two approaches. Nevertheless, this highlights the need to ensure spectrometer geometry still allows the full intensity of the signal to be captured within the spectral region of interest, as a sharp drop off such as this would certainly lead to different conclusions from the actual conditions. Furthermore, the relation in position of the source, crystal, and camera are important to the specific shape of the spectrometer IF.
@@ III.4 Temperature via Detailed Balance
Figure 12: Comparison between a simulated experimental spectrum of CH at 83.1 eV using the 40 $\mu$m crystal SIF (black solid), and a fit to this spectrum but with a 4 eV FWHM Gaussian as the estimated SIF (red dashed); a comparison of these two SIFs is shown in the inset. Also plotted is the original DSF convolved with the 4 eV Gaussian (green dot-dashed), and the original spectrum with Gaussian noise [cf. Eq. (12)] with standard deviation $s=0.05$ applied.
@@
Finally, we consider the effect of ignoring the asymmetry in the crystal IF on plasma conditions inferred from a XRTS spectrum in experiment. In Fig. 12, a simulated XRTS spectrum for a CH plasma at a temperature $T=83.1$ eV, mean ionization state $Z=3.51$, and mass density $\rho=1.2$ g/cm3. The DSF is calculated using the multi-component scattering simulation (MCSS) code Chapman (2015), with the convolution of the DSF with the SIF performed externally to produce the final XRTS spectrum. The DSF has been probed with a monochromated 8.5 keV beam and measured at a scattering angle of 17 on the 40 $\mu$m crystal. In this case, we consider the effect of the SIF as a convolution to isolate the effect of the asymmetry in a single SIF, noting that the relative differences plotted in Fig. 10 indicates the full SIF can show further asymmetry between the upshifted and downshifted sides of the spectrum versus a simple convolution.
@@
The fact that all those plots give reasonable looking fits to the noisy data is not entirely surprising – the forward modelling of XRTS spectra is a notoriously unstable inverse problem when the spectrum has noise Kasim et al. (2019). However, recently laser systems at XFEL facilities have been implemented with repetition rates comparable to the XFEL ($\sim 1-10$ Hz), such as ReLaX Laso Garcia et al. (2021) and DiPOLE-100X Mason et al. (2018) at the HED endstation at EuXFEL. It is therefore now possible to rapidly collect 1,000s–10,000s of shots of warm and hot dense matter systems produced using the laser and probed with the bright FEL beam. The photon collection statistics in these setups will therefore greatly improve, and the noise diminished. As such, these subtle differences in the fitting may become more apparent, particularly as techniques to reliably remove the background to the spectra become more developed. It is anticipated the accurate determination of the SIF will become increasingly important to the accurate forward fitting of the XRTS spectrum.
@@ IV Conclusion
We have presented detailed calculations of the IF of mosaic crystals. Previously, the weigh up of using mosaic crystals has been between their very high integrated reflectivities versus their relatively poor resolution. However, we have demonstrated the IFs of these crystals are very asymmetric to higher energies in the geometries considered here. If this asymmetry is not accounted for, it has a strong impact on the inferred plasma conditions even when the IF may appear to look reasonably symmetric in a spectrum. Notably, we find the temperature inferred via the detailed balance relation for an XRTS spectrum can be substantially higher when using a symmetric function, as the DSF needs to compensate for the true IF boosting the upshifted side of the spectrum. This applies in both the forward modelling and ITCF analysis.
We also find that the instrument function is not a convolution, but is instead a kernel that depends on the photon energy in a non-trivial manner. Furthermore, the specific shape of the IF depends on the geometry of the set up, which results in even greater deviations and some notable features within the IF that vary across the camera. Nevertheless, for a sufficiently narrow spectral range the kernel integration can be approximated as a convolution, which greatly simplifies forward model fitting to XRTS spectra, and allows the model-free ITCF method to be employed with high accuracy. However, the applicable energy range is still relatively narrow and will depend on the spectrum being measured. Furthermore, the geometry dependence of the spectrometer IF suggests that while detailed studies of the spectrometer components are very useful for understanding the behaviour of the spectrometer, a measurement of the specific IF for a given experiment is still important.
Here, we have only considered the effect of the asymmetry of the SIF on the inferred temperature. Further work is required to investigate the influence of the full SIF on the inference of other plasma properties inferred via XRTS, such as the frequency moments of the DSF Dornheim et al. (2023c). Nevertheless, we expect the choice of SIF model will have an effect on properties inferred.
@@ Acknowledgments
We acknowledge the European XFEL in Schenefeld, Germany, for provision of X-ray free-electron laser beamtime at the Scientific Instrument HED (High Energy Density Science) under proposal numbers 3777 and 5690 and would like to thank the staff for their assistance. The authors are grateful to the HIBEF user consortium for the provision of instrumentation and staff that enabled these experiments.
@@
The authors have no conflicts of interest to disclose. T.G., T.D., H.B. performed the analysis, with input from Z.M., J.V.. T.G. developed the code for the model, with contributions form L.B.F. and M.J.M.. T.G., H.B., K.A., C.B., V.B., E.B., D.B., A.C., A.D., S.G., N.J.H., M.-L.H, P.H., H.H., O.S.H., Z.K., A.L., B.L., J.L., M.M., W.M., M.M. Z.A.M., M.N., J.-P.N, P.N., A.P., C.Q., L.R., J.R., T.T., L.W., U.Z., D.K., T.R.P., T.D. were involved in the collection of the experimental data. T.R.P. was the main proposer and principal investigator of experiment p3777. D.K. was the main proposer and principal investigator of experiment p5690. T.G. wrote substantial parts of the manuscript, with contribution from T.D.. All authors were provided the opportunity to review the manuscript.

### v2

@@ III.3 Source-and-Instrument Function as a Convolution
In Fig. 10, we compare calculations of simulated XRTS spectra using the convolution and for a full kernel calculation. The simulated spectra are for a CH plasma at a temperature $T=83.1$ eV, mean ionization state $Z=3.51$, and mass density $\rho=1.2$ g/cm3 calculated using the multi-component scattering simulation (MCSS) code Chapman (2015), with the DSF probed with a monochromated 8.5 keV beam and measured at scattering angles of 17 and 166. In the case of the convolution calculation, the convolution of the DSF with the SIF performed externally to produce the final XRTS spectrum. To estimate the energy dependence of the mosaicity and IRC, we use linear fits to the data reported Ref. Gerlach et al. (2015), maintaining the gradients but shifting the values to match the mosaic spread and IRC widths we report in Table 1 (b).
@@
In backwards scattering, the differences between the full kernel calculation and convolution become more apparent, with the relative difference increasing further from the quasi-elastic feature. This is because, compared to forward scattering, there is substantial signal at energy losses far from the quasi-elastic feature in backwards scattering which will contribute to the differences in shape across the spectrum. Additionally, there is a notable drop in intensity above 8.6 keV – this is because the crystal position was set only to measure up to $\sim 100$ eV above the elastic (cf. Fig. 7 (b)), so it is perhaps not surprising to see substantial differences between the two approaches. Nevertheless, this highlights the need to ensure spectrometer geometry still allows the full intensity of the signal to be captured within the spectral region of interest, as a sharp drop off such as this would certainly lead to different conclusions from the actual conditions. Furthermore, the relation in position of the source, crystal, and detector are important to the specific shape of the spectrometer IF.
@@ III.4 Temperature via Detailed Balance
Figure 12: Comparison between a simulated experimental spectrum (red-dashed curve in Fig. 10 (a)) of CH at 83.1 eV using the 40 $\mu$m crystal SIF (black solid), and a fit to this spectrum but with a 4 eV FWHM Gaussian as the estimated SIF (red dashed); a comparison of these two SIFs is shown in the inset. Also plotted are the DSF for the original black solid curve convolved with the 4 eV Gaussian (green dot-dashed), and the original solid black curve with Gaussian noise [cf. Eq. (13)] with standard deviation $s=0.05$ applied (blue dotted).
@@
Finally, we consider the effect of ignoring the asymmetry in the crystal IF on plasma conditions inferred from a XRTS spectrum in experiment. In Fig. 12, the simulated spectra is the same as the red-dashed curve in Fig. 10 (a). We consider here the effect of the SIF as a convolution to isolate the effect of the asymmetry in a single SIF, noting that the relative differences plotted in Fig. 10 indicates the full SIF can show further asymmetry between the upshifted and downshifted sides of the spectrum versus a simple convolution.
@@
The fact that all those plots give reasonable looking fits to the noisy data is not entirely surprising – the forward modelling of XRTS spectra is a notoriously unstable inverse problem when the spectrum has noise Kasim et al. (2019). However, recently laser systems at XFEL facilities have been implemented with repetition rates comparable to the XFEL ($\sim 1-10$ Hz), such as ReLaX Laso Garcia et al. (2021) and DiPOLE-100X Mason et al. (2018) at the HED instrument at EuXFEL. It is therefore now possible to rapidly collect 1,000s–10,000s of shots of warm and hot dense matter systems produced using the laser and probed with the bright FEL beam. The photon collection statistics in these setups will therefore greatly improve, and the noise diminish. As such, these subtle differences in the fitting may become more apparent, particularly as techniques to reliably remove the background to the spectra become more developed. It is anticipated the accurate determination of the SIF will become increasingly important to the accurate forward fitting of the XRTS spectrum.
@@ IV Conclusion
We have presented detailed calculations of the IF of mosaic crystals. Previously, the benefit of using mosaic crystals has been between their very high integrated reflectivities versus their relatively poor resolution. However, we have demonstrated the IFs of these crystals are very asymmetric to higher energies in the geometries considered here. The asymmetry arises due to the depth broadening and the mosaic distribution of the crystallites, both of which, when isolated, only result in broadening towards higher energies. Broadening towards lower energies is facilitated by the intrinsic rocking curve of the crystallites. If this asymmetry is not accounted for, it has a strong impact on the inferred plasma conditions even when the IF may appear to look reasonably symmetric in a spectrum. Notably, we find the temperature inferred via the detailed balance relation for an XRTS spectrum can be substantially higher when using a symmetric function, as the DSF needs to compensate for the true IF boosting the upshifted side of the spectrum. This applies in both the forward modelling and ITCF analysis.
We also find that the instrument function is not a convolution, but is instead a kernel that depends on the photon energy in a non-trivial manner. Furthermore, the specific shape of the IF depends on the geometry of the set up, which results in even greater deviations and some notable features within the IF that vary across the detector. Nevertheless, for a sufficiently narrow spectral range the kernel integration can be approximated as a convolution, which greatly simplifies forward model fitting to XRTS spectra, and allows the model-free ITCF method to be employed with high accuracy. However, the applicable energy range is still relatively narrow and will depend on the spectrum being measured. Furthermore, the geometry dependence of the spectrometer IF suggests that while detailed studies of the spectrometer components are very useful for understanding the behaviour of the spectrometer, a measurement of the specific IF for a given experiment is still important.
Here, we have only considered the effect of the asymmetry of the SIF on the inferred temperature. Further work is required to investigate the influence of the full SIF on the inference of other plasma properties inferred via XRTS, such as the frequency moments of the DSF Dornheim et al. (2023c), the plasma density, and ionization. Nevertheless, we expect the choice of SIF model will have an effect on properties inferred.
@@ Acknowledgments
We acknowledge the European XFEL in Schenefeld, Germany, for provision of X-ray free-electron laser beamtime at the Scientific Instrument HED (High Energy Density Science) under proposal numbers 3777 and 5690 and would like to thank the staff for their assistance. The authors are grateful to the HIBEF user consortium for the provision of instrumentation and staff that enabled these experiments. The original datasets can be found here and are available upon reasonable request: doi:10.22003/XFEL.EU-DATA-003777-00, doi:10.22003/XFEL.EU-DATA-005690-00.
@@
The authors have no conflicts of interest to disclose. T.G., T.D., H.B. performed the analysis, with input from Z.M., J.V.. T.G. developed the code for the model, with contributions form L.B.F. and M.J.M.. T.G., H.B., K.A., C.B., V.B., E.B., D.B., A.C., A.D., S.G., N.J.H., M.-L.H, P.H., H.H., O.S.H., Z.K., A.L., B.L., J.L., M.M., W.M., M.M. Z.A.M., M.N., J.-P.N, P.N., A.P., C.Q., L.R., J.R., T.T., L.W., U.Z., D.K., T.R.P., T.D. were involved in the collection of the experimental data. T.R.P. was the main proposer and principal investigator of experiment proposal 3777. D.K. was the main proposer and principal investigator of experiment proposal 5690. T.G. wrote substantial parts of the manuscript, with contribution from T.D. and T.R.P. All authors were provided the opportunity to review the manuscript.